# Tuning-Free Adaptive Style Incorporation for Structure-Consistent Text-Driven Style Transfer

### v1

@@ Abstract
In this work, we target the task of text-driven style transfer in the context of text-to-image (T2I) diffusion models. The main challenge is consistent structure preservation while enabling effective style transfer effects. The past approaches in this field directly concatenate the content and style prompts for a prompt-level style injection, leading to unavoidable structure distortions. In this work, we propose a novel solution to the text-driven style transfer task, namely, Adaptive Style Incorporation (ASI), to achieve fine-grained feature-level style incorporation. It consists of the Siamese Cross-Attention (SiCA) to decouple the single-track cross-attention to a dual-track structure to obtain separate content and style features, and the Adaptive Content-Style Blending (AdaBlending) module to couple the content and style information from a structure-consistent manner. Experimentally, our method exhibits much better performance in both structure preservation and stylized effects.
Keywords:
Style transfer Diffusion model
00footnotetext: $*{}^{*}$qual contribution.
${}^{\dagger}$orrespoding author.
@@
1 Introduction
With the increasing prominence of text-to-image (T2I) [37, 35, 36] models featuring billion-parameter architectures trained on extensive text-image datasets, the prospect of translating our imaginative ideas into high-quality images through textual input becomes readily apparent. While these foundation models have ushered in a new era of creative expression, their powerful generation capability comes at the expense of limited controllability. Due to the text-image misalignment problem in the T2I models, even minor modifications to the textual prompt may result in unforeseen structure distortion in the output image. This poses a significant obstacle for artists and creators seeking precise artistic editing to their work.
In this study, we aim to investigate text-driven style transfer (see Fig. 1) within the context of T2I diffusion models. The major challenge in this field lies in preserving the original structure while producing effective stylized results. To resolve the aforementioned issue, numerous efforts have been made to incorporate additional conditional controls, such as structural guidance signals [49, 52] and editing masks [12, 2]. To be specific, BlendDiffusion [4, 2] and SpaText [3] train diffusion models that support a user-provided mask as input to achieve regional generation and editing. ControlNet [52] and Make-A-Scene [12] finetune diffusion models conditioned on the segmentation masks to ensure invariance of semantics in the editing images. However, these methods fall short of achieving fine-grained control over the structural layout and require extensive training on large-scale text-image pairs, thereby hindering intuitive and easily deployable text-driven style transfer.
@@ 1 Introduction
However, the aforementioned approaches choose to directly concatenate the content prompt (representing the source image) and style prompt, in order to achieve prompt-level style injection. Nevertheless, due to the well-known text-image misalignment issue [23] in T2I models, the style prompt often contains non-style information, leading to unavoidable structure distortions and semantic drift in the stylized image for P2P and PnP.
Unlike previous studies that commonly inject styles on the coarse-grained prompt level, we propose Adaptive Style Incorporation (ASI) to incorporate the style information on the fine-grained feature level. In ASI, we first present the Siamese Cross-Attention (SiCA) that decouples the common single-track cross attention into a dual-track architecture, which utilizes the shared Query features to simultaneously query two sets of Key and Value features from the content prompt and style prompt individually, as depicted in Figure. 2 (b). Subsequently, we introduce a tuning-free Adaptive Content-Style Blending (AdaBlending) module to implement mask-guided style incorporation from a structure-consistent perspective. Thanks to the proposed ASI mechanism, our feature-level fine-grained style incorporation can maintain better structure consistency for the style transfer task.
@@
  • We propose the Siamese Cross-Attention (SiCA) and Adaptive Content-Style Blending (AdaBlending) modules to form the complete ASI mechanism. It requires no training/fine-tuning but rather leverages a pre-trained text-to-image diffusion model.
  • Experimental results have shown the effectiveness of our proposed ASI in both maintaining the structure-consistent layout and effectively embedding style attributes, whether applied to real images or generated images.
2 Related Work
2.1 Diffusion Models
The Diffusion Model [40] was initially employed to mitigate Gaussian noise applied continuously to training images and can be conceived as a series of denoising autoencoders. The advent of Denoising Diffusion Probabilistic Models (DDPM) [21] propelled the utilization of diffusion models for image generation into the mainstream. Subsequent advancements, exemplified by Denoising Diffusion Implicit Models (DDIM) [41], substantially accelerated sampling speeds while effecting directed generation from random noise to samples, facilitating subsequent research in image editing. Dhariwal and Nichol first introduced Classifier Guidance Diffusion [11], substantiating the superiority of diffusion models over GANs in image generation tasks. Classifier-Free Diffusion Guidance [22] then achieved a balance between sample quality and diversity during image generation. Building upon prior work, GLIDE [34] initiated using text as a conditional guide for image generation in a classifier-free manner. DALLE-2 [35], building on the foundation laid by GLIDE, incorporated CLIP as a guiding factor, bringing diffusion models into the public eye. Subsequently, the field of image generation based on diffusion models experienced a surge, with projects such as Imagen [37] and Stable Diffusion [36] emerging, prompting the gradual application of diffusion models in a broader range of tasks, including object detection [7], style transfer [46], image editing [29, 5, 8] and semantic segmentation [51]. With the release of ControlNet [52], the potential of diffusion models in various other domains is being further explored [10].
2.2 Style Transfer
@@ 3 Preliminaries
[…]
where $\alpha_{t}=1-\beta_{t}$ and $\hat{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$.
@@
[…]
@@
Figure 2: Comparison between prompt-level coarse-grained style injection with our feature-level fine-grained style incorporation. (a) Due to the well-known issue of text-image misalignment in T2I models, directly concatenating the content prompt with the style prompt (i.e., prompt-level coarse-grained style injection) will introduce non-style information into the target image, resulting in unavoidable structural and semantic drift in the target image, such as the hair of the boy and the pattern on his clothes. (b) Compared to prompt-level coarse-grained style injection, we propose Adaptive Style Incorporation (ASI) to convert the existing cross-attention module to achieve feature-level fine-grained style incorporation. ASI consists of the proposed Siamese Cross-Attention (SiCA) and Adaptive Content-Style Blending (AdaBlending) modules, which collaborate to preserve image structure while facilitating effective style transfer.
4 Method
@@
As discussed in Sec. 1, existing methods [19, 43, 32] commonly directly concatenate the content prompt and style prompt (e.g., a boy in anime artwork) and transform to key-value feature pair to achieve prompt-level style injection. We question this coarse-grained form of prompt-level style injection. Our motivation stems from the fact that if we do not reuse the spatial features of the reference diffusion process, and simply add a style token to the content prompt, the text-image misalignment problem in the T2I model will often result in unpredictable structural drift in the stylized image $I_{s}$. In this sense, if we explicitly parse style information that does not disrupt the structure of content features and incorporate it into content features, we can achieve more controllable results.
For this purpose, we propose the Adaptive Style Incorporation (ASI) to convert the existing cross-attention module to achieve structure-consistent feature-level fine-grained style transfer. The overall architecture of the proposed ASI is depicted in Fig. 2, which is composed of a Siamese Cross-Attention (SiCA) strategy and Adaptive Content-Style Blending (AdaBlending) module. We first introduce SiCA which employs a dual-track architecture to obtain independent content and style feature-level information (Sec. 4.1). Then, we propose the AdaBlending module to implement mask-guided fine-grained style incorporation from a structure-consistent perspective (Sec. 4.2). The proposed method can be applied to style transfer for both real and generated images without requiring training or fine-tuning, achieving a highly structurally consistent stylized image with the content image.
4.1 Siamese Cross-Attention
[…]
@@ 4.1 Siamese Cross-Attention
This design aims to empower the query features, which signify the image structure, to independently attend to both the style and content prompt embeddings. Compared to prompt-level coarse-grained one-shot style injection, we can initially extract the style information from $F_{s}$ that has minimal structural impact on $F_{c}$ and adaptively incorporate it into $F_{c}$, achieving feature-level fine-grained style incorporation. We will detail it in the next section.
4.2 Adaptive Content-Style Blending
We propose a tuning-free Adaptive Content-Style Blending (AdaBlending) module to implement mask-guided fine-grained style incorporation from a structure-consistent perspective. AdaBlending first utilizes two parallel mask extractors to parse out the content-style blending region from the attention-head (Sec. 4.2.1) and spatial levels (Sec. 4.2.2), then employs the mask-guided content-style blending module (Sec. 4.2.3) to effectively stylize the target image while maintaining its structural consistency with the source content image. The overall architecture of the proposed AdaBlending is shown in Fig. 3.
Figure 3: Overall architecture of the proposed Adaptive Content-Style Blending (AdaBlending) module. AdaBlending achieves mask-guided fine-grained style incorporation through the proposed attention-head level and spatial level mask extractors, followed by the content-style blending operation in the structure-consistent manner.
4.2.1 Mask Extractor on Attention-Head Level
Many pioneering works  [24, 20, 44, 33, 57, 31, 30, 25, 53] have demonstrated that the spatial distribution of features effectively represents the style of an image. Therefore, if an attention head does not exhibit significant distribution differences in $F_{s}$ and $F_{c}$, we consider incorporating style information into this attention head is ineffective and disrupts the image structure of the target image.
[…]
@@ 4.2.1 Mask Extractor on Attention-Head Level
where $Cov_{s}$ and $Cov_{c}$ denote the spatial level covariance of style and content features, respectively. ${\|\cdot\|}^{2}_{F}$ represents the squared matrix Frobenius norm, 1 is a vector with all elements equal to 1, $m$, $d$, $i$ are flattened spatial dimension, channel dimension, and attention-head index, respectively. Then, we can obtain the attention-head mask $M_{head}\in\mathbb{R}^{h\times m\times d}$ as follows,
$M_{head}^{i}=\begin{cases}\textbf{1},\quad&\text{if}~{}~{}~{}~{}\mathop{\arg\max}_{i}\ \ (\ell^{i},n)\\ 0,\quad&\text{otherwise}\\ \end{cases}\,$ (15)
@@
where $\mathop{\arg\max}_{i}~{}(\ell^{i},n)$ denotes the return of the index of top-$n$ attention heads with the greatest covariance distances, and $n$ is set to 6 for Stable Diffusion v1 [36]. $h$ denotes the number of heads of multi-head cross-attention.
4.2.2 Mask Extractor on Spatial Level
@@
We further consider how to maintain structure consistency from the spatial level. Many works [56, 38, 28, 27] have demonstrated that regions with high activation correspond to crucial structural information within images. Therefore, we propose a mask extractor on the spatial level to maintain core structure information of $F_{c}$. The extractor selects criteria points of peak regions using the global max pooling of $F_{c}$, and the criteria points are denoted as $P\in\mathbb{R}^{h\times 1\times d}$. We adopt a non-learnable setting for the extractor and define regions in $F_{c}$ exceeding $\tau=\alpha\cdot P$ as crucial structural regions. $\alpha$ is a constant value and set to $0.7$. To preserve the features of these highly activated regions, we formulate the spatial mask $M_{spatial}\in\mathbb{R}^{h\times m\times d}$ at each attention-head as follows,
$M_{spatial}^{i}=\begin{cases}0,\quad&\text{if}~{}~{}~{}~{}F_{c}^{i}>\tau\\ \textbf{1},\quad&\text{otherwise}\\ \end{cases}\,$ (16)
4.2.3 Mask-Guided Content-Style Blending
With the above masks, we restrict the content features $F_{c}$ to incorporate style features $F_{s}$ from a structural-consistent perspective. We first perform logical OR operations on two sets of masks $M_{head}$ and $M_{spatial}$ to obtain the final mask $M_{all}$. Then, we perform the mask-guided blending between the structured style features powered by adaptive instance normalization operation [24], and content features to obtain the fused fine-grained stylized result as AdaBlending’s output $F_{out}$,
$F^{i}_{out}=(\sigma(F_{s}^{i})(\frac{F_{c}^{i}-\mu(F_{c}^{i})}{\sigma(F_{c}^{i})})+\mu(F_{s}^{i}))\odot M_{all}^{i}+F_{c}^{i}\odot(\textbf{1}-M_{all}^{i}),$ (17)
where $\odot$ denotes the inner product operation.
5 Experiments
@@
In this section, we perform both qualitative and quantitative evaluations across three applications: style transfer for real-world images, visual enhancement for real-world images, and style transfer for generated images. We compare our method against state-of-the-art (SOTA) tuning-free text-guided style transfer method based on diffusion models, including PnP [43], P2P [19], and SDEdit [32], across several common style transfer tasks. Finally, we ablate each component of our method and show its effectiveness. We evaluate our method using Stable Diffusion v1 architecture [36]. Please refer to the Appendix for the implementation details of our method, quantitative evaluation results, and more quantitative results.
5.1 Qualitative Comparison
5.1.1 Style Transfer for Real-World Images
Figure 4: Comparisons with different methods on style transfer for real-world images. Sample results are shown for each of the two benchmarks: ImageNet-R-TI2I and Wild-TI2I. From left to right: the source image and text prompt, our results, P2P [19], PnP [43], SDEdit [32] with two different denoising strengths. (Zoom in for best view.)
To verify the effectiveness and generalization of ASI, we show various style transfer results applied by our approach on Wild-TI2I [43] and ImageNet-R-TI2I [18, 43] datasets, which are collected from real-world. For each result, we show images before and after stylized. As shown in Fig. 4, we compare our method against SOTA text-guided style image editing models based on diffusion models, including PnP [43], P2P [19], and SDEdit [32]. Since our method aims to incorporate style information at the fine-grained feature level, compared to prompt-level style injection, it allows us to stylize in the smaller local region of images, thereby reducing the distortion of image structure. Although SDEdit is simple but effective, it requires manual adjustment of the trade-off between the two. PNP [43] focuses more on the structural consistency of stylized images but lacks effective stylization in some cases (e.g. Row. 3 and 5). While P2P [19] is the opposite, effective stylization comes at the expense of structural and instance deviation (e.g. stone in the Row. 3). As shown in Col. 2 of Fig. 4, our method effectively stylizes images while maintaining structural consistency with the source image, achieving a good balance between stylization and structural preservation compared to other SOTA methods.
Figure 5: Comparisons with different methods on visual enhancement for real-world images. From left to right: the source image, visual enhancement prompts, our results, P2P [19], PnP [43], SDEdit [32]. (Zoom in for best view.)
@@ 5.1.1 Style Transfer for Real-World Images
Figure 6: Comparisons with different methods on style transfer for generated images. From left to right: the source image and text prompt, our results, P2P [19], PnP [43], SDEdit [32] with two different denoising strengths. (Zoom in for best view.)
5.1.2 Visual Enhancement for Real-World Images

### v2

@@ Abstract
Text-driven style transfer methods leveraging diffusion models have shown impressive creativity, yet they still face challenges in maintaining consistent structure and content preservation. Existing methods often directly concatenate the content and style prompts for a prompt-level style injection. However, this coarse-grained style injection strategy inevitably leads to structural deviations in the stylized images. This poses a significant obstacle for professional artists and creators seeking precise artistic editing. In this work, we strive to attain a harmonious balance between content preservation and style transformation. We propose Adaptive Style Incorporation (ASI), to achieve fine-grained feature-level style incorporation. It consists of the Siamese Cross-Attention (SiCA) to decouple the single-track cross-attention to a dual-track structure to obtain separate content and style features, and the Adaptive Content-Style Blending (AdaBlending) module to couple the content and style information from a structure-consistent manner. Experimentally, our method exhibits much better performance in both structure preservation and stylized effects.
@@
I Introduction
With the increasing prominence of text-to-image (T2I) [41, 38, 40] models featuring billion-parameter architectures trained on extensive text-image datasets, the prospect of translating our imaginative ideas into high-quality images through language becomes readily apparent. One thrilling application is the image style transfer via textual inputs. However, due to the text-image misalignment problem in the T2I models, even minor modifications to the textual prompt may result in unforeseen structure distortion in the output image. This poses a significant obstacle for professional creators who aim for artistic renderings while precisely preserving the semantics, structure, and context of the source image.
In this work, we strive to attain a harmonious balance between content preservation and style transformation. To achieve this objective, numerous efforts have been made to incorporate additional conditional controls, such as structural guidance signals [57, 60] and editing masks [14, 2]. To be specific, BlendDiffusion [4, 2] and SpaText [3] train diffusion models that support a user-provided mask as input to achieve regional generation and editing. ControlNet [60] and Make-A-Scene [14] finetune diffusion models conditioned on the segmentation masks to ensure invariance of semantics in the editing images. However, these methods require extensive training on large-scale text-image pairs and fall short of achieving fine-grained control over the structural layout, thereby hindering intuitive and easily deployable text-driven style transfer.
@@ I Introduction
However, the aforementioned approaches choose to directly concatenate the content prompt (representing the source image) and style prompt, to achieve prompt-level style injection. Nevertheless, due to the well-known text-image misalignment issue [24] in T2I models, the style prompt often contains non-style information, leading to unavoidable structure distortions and semantic drift in the stylized image for P2P and PnP.
Unlike previous studies that commonly inject styles on the coarse-grained prompt level, we propose Adaptive Style Incorporation (ASI) to incorporate the style information on the fine-grained feature level. In ASI, we first present the Siamese Cross-Attention (SiCA) that decouples the common single-track cross-attention into a dual-track architecture, which utilizes the shared Query features to simultaneously query two sets of Key and Value features from the content prompt and style prompt individually, as depicted in Fig. 2 (b). Subsequently, we introduce a tuning-free Adaptive Content-Style Blending (AdaBlending) module to parse style features that do not disrupt the structure of content features and incorporate them into the content features. Thanks to the fine-grained style incorporation, the ASI mechanism can preserve better structure consistency between content and stylized images.
@@
  • We propose the Siamese Cross-Attention (SiCA) and Adaptive Content-Style Blending (AdaBlending) modules to form the complete ASI mechanism. It requires no training or fine-tuning but rather leverages a pre-trained text-to-image diffusion model.
  • Experimental results show that our proposed method effectively embeds style attributes while preserving the structural layout of the content image, regardless of whether it is applied to real or generated images.
II Related Work
II-A Diffusion Models
The Diffusion Model [44] was initially employed to mitigate Gaussian noise applied continuously to training images and can be conceived as a series of denoising autoencoders. The advent of Denoising Diffusion Probabilistic Models [22] propelled the utilization of diffusion models for image generation into the mainstream. Subsequent advancements, exemplified by Denoising Diffusion Implicit Models [45], substantially accelerated sampling speeds while effecting directed generation from random noise to samples, facilitating subsequent research in image editing. Dhariwal and Nichol first introduced Classifier Guidance Diffusion [13], substantiating the superiority of diffusion models over GANs in image generation tasks. Classifier-Free Diffusion Guidance [23] then achieved a balance between sample quality and diversity during image generation. Building upon prior work, GLIDE [35] initiated using text as a conditional guide for image generation in a classifier-free manner. DALLE-2 [38], building on the foundation laid by GLIDE, incorporated CLIP as a guiding factor, bringing diffusion models into the public eye. Subsequently, the field of image generation based on diffusion models experienced a surge, with projects such as Imagen [41] and Stable Diffusion [40] emerging, prompting the gradual application of diffusion models in a broader range of tasks, including object detection [8], style transfer [53], image editing [30, 5, 9] and semantic segmentation [59]. With the release of ControlNet [60] and IP-Adapter [57], the use of more diverse and richer conditions for image editing and customized image generation is gaining widespread exploration and application.
II-B Style Transfer
@@ III Preliminaries
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$.
@@
DDIM [45] introduces an alternative non-Markovian process, allowing the denoising process to be different. Given timestep $x_{t}$ and $\epsilon_{\theta}(x_{t},t)$ at timestep $t$, DDIM can predict $x_{0}=f_{\theta}(x_{t},t)$ and $x_{t-1}$ as follows,
@@
Figure 2: (a) Due to the text-image misalignment in T2I models, directly concatenating the content prompt with the style prompt (i.e., prompt-level coarse-grained style injection) will introduce non-style information into the style transfer process, resulting in unavoidable structural and semantic drift in the stylized image, such as the hair of the boy and the pattern on his clothes. (b) We propose Adaptive Style Incorporation (ASI), which consists of siamese cross-attention and adaptive content style blending modules. ASI explicitly parses style information that does not disrupt the structure of content features and incorporates it into the content features to achieve feature-level fine-grained style incorporation.
Figure 3: Overall architecture of our proposed Adaptive Content-Style Blending (AdaBlending) module. AdaBlending achieves mask-guided fine-grained style incorporation through the proposed attention-head level and spatial level mask extractors, followed by the content-style blending operation in the structure-consistent manner.
IV Proposed Method
@@
As discussed in Sec. I, existing methods [20, 50, 33] commonly directly concatenate the content prompt and style prompt (e.g., a boy in anime artwork) and transform to key-value feature pair to achieve prompt-level style injection. We question this coarse-grained form of prompt-level style injection. Our motivation stems from the fact that if we do not reuse the spatial features of the reference diffusion process, and simply add a style token to the content prompt, the text-image misalignment problem in the T2I model will often result in unpredictable structural drift in the stylized image $I_{s}$. In this sense, we gain greater control over the results if we explicitly parse style information that does not disrupt the structure of content features and incorporate it into content features.
For this purpose, we propose the Adaptive Style Incorporation (ASI) to convert the existing cross-attention module to achieve structure-consistent feature-level fine-grained style transfer. The overall architecture of the proposed ASI is depicted in Fig. 2, which is composed of a Siamese Cross-Attention (SiCA) strategy and Adaptive Content-Style Blending (AdaBlending) module. We first introduce SiCA which employs a dual-track architecture to obtain independent content and style feature-level information (Sec. IV-A). Then, we propose the AdaBlending module to implement mask-guided fine-grained style incorporation from a structure-consistent perspective (Sec. IV-B). The proposed method can be applied to style transfer for both real and generated images without requiring training or fine-tuning, achieving a highly structurally consistent stylized image with the content image.
IV-A Siamese Cross-Attention
As depicted in Fig. 2 (b), we first present a Siamese Cross-Attention (SiCA) that decouples the common single-track cross-attention into a dual-track architecture, where the two branches of the architecture share the parameters of the replaced cross-attention. We call these two branches the “style branch” and “content branch”. In addition, we formalize the output features $F_{s}$ and $F_{c}$ as follows,
@@ IV-A Siamese Cross-Attention
This design aims to empower the query features, which signify the image structure, to independently attend to both the style and content prompt embeddings. Compared to prompt-level coarse-grained one-shot style injection, we can initially extract the style information from $F_{s}$ with minimal structural impact on $F_{c}$ and adaptively incorporate it into $F_{c}$, achieving feature-level fine-grained style incorporation. We will detail it in the next section.
IV-B Adaptive Content-Style Blending
We propose a tuning-free Adaptive Content-Style Blending (AdaBlending) module to implement mask-guided fine-grained style incorporation from a structure-consistent perspective. The overall architecture of the AdaBlending is shown in Fig. 3. AdaBlending first utilizes two parallel mask extractors to parse out the style incorporation region, which is composed of an attention-head mask $M_{head}$ and spatial mask $M_{spatial}$. The attention-head mask $M_{head}$ ensures that the style injection performed on each attention head is effective, while the spatial mask $M_{spatial}$ prevents style injection on the high-frequency edge information of the content image, thereby preserving the integrity of the content object.
With these masks, AdaBlending restricts the content features $F_{c}$ to incorporate style features $F_{s}$ from a structurally consistent perspective. We first perform logical OR operations on two sets of masks $M_{head}$ and $M_{spatial}$ to obtain the final mask $M_{all}$. Then, we perform the mask-guided blending between the structured style features powered by adaptive instance normalization operation [26], and content features to obtain the fused fine-grained stylized result as AdaBlending’s output $F_{out}$,
$F^{i}_{out}=$ $(\sigma(F_{s}^{i})(\frac{F_{c}^{i}-\mu(F_{c}^{i})}{\sigma(F_{c}^{i})})+\mu(F_{s}^{i}))\odot M_{all}^{i}$ (12)
$+F_{c}^{i}\odot(\textbf{1}-M_{all}^{i}),$
where $\odot$ denotes the inner product operation. the $\mu(\cdot)$ and $\sigma(\cdot)$ represent functions for taking the standard deviation and the mean of features, respectively. The next two sections will detail how to generate the $M_{head}$ and $M_{spatial}$.
IV-B1 Mask Extractor on Attention-Head Level
Many pioneering works  [26, 21, 51, 34, 64, 32, 31, 27, 61, 11] have demonstrated that the spatial distribution of features effectively represents the style of an image. Therefore, if an attention head does not exhibit significant distribution differences in $F_{s}$ and $F_{c}$, we consider incorporating style information into this attention head ineffective and disrupts the stylized image’s image structure.
We measure the distribution difference [47] between each pair of attention heads in $F_{s}$ and $F_{c}$ by calculating the distance of covariance $\ell^{i}$ as follows,
@@ IV-B1 Mask Extractor on Attention-Head Level
where $Cov_{s}$ and $Cov_{c}$ denote the spatial level covariance of style and content features, respectively. ${\|\cdot\|}^{2}_{F}$ represents the squared matrix Frobenius norm, 1 is a vector with all elements equal to 1, $m$, $d$, $i$ are flattened spatial dimension, channel dimension, and attention-head index, respectively. Then, we can obtain the attention-head mask $M_{head}\in\mathbb{R}^{h\times m\times 1}$ as follows,
[…]
@@
where $\mathop{\arg\max}_{i}~(\ell^{i},n)$ denotes the return of the index of top-$n$ attention heads with the greatest covariance distances. $n$ is set to 6 for visual enhancement and to 7 for style transfer in the Stable Diffusion V1 architecture. $h$ denotes the number of heads of multi-head cross-attention.
Figure 4: Visualization of the top three leading components of average cross-attention features of the U-Net.
IV-B2 Mask Extractor on Spatial Level
We further consider how to maintain structure consistency from the spatial level. Inspired by previous work [20, 50, 6, 48], the leading principal components of the attention features implicitly encode the image core structure information. As shown in Fig. 4, the highly activated positions basic correspond to the image’s high-frequency low-level edge information.
@@ IV-B2 Mask Extractor on Spatial Level
Therefore, we propose a spatial-level mask extractor to achieve style incorporation exclusively in non-structural regions. The extractor selects criteria points of peak regions using the global max pooling of $F_{c}$’s first principal component, and the criteria points are denoted as $P\in\mathbb{R}^{h\times m\times 1}$. We adopt a non-learnable setting for the extractor and define regions in $F_{c}$’s first principal component exceeding $\tau=\alpha\cdot P$ as crucial structural regions. $\alpha$ is a constant value and set to $0.7$.
We formulate the spatial mask $M_{spatial}\in\mathbb{R}^{h\times m\times 1}$ at each attention-head as follows,
[…]
Figure 5: Sample results of our method for transferring both photography and artistic styles. Zoom in for the best view.
V Experiments
@@
In this section, we conduct both qualitative and quantitative evaluations of style transfer and visual enhancement tasks. These evaluations are performed on both real-world images and images generated from given prompts using the Stable Diffusion model. We focus our comparisons on SOTA tuning-free text-driven style transfer method based on diffusion models, including FreeDiff [55], PnP [50], P2P [20], and SDEdit [33]. In addition, we further qualitatively compare our method with other methods that have different experimental settings, including AdaIN [26], AdaAttn [32], Z* [12], DiffStyler [25], ClipStyler [29], IP-Adapter [57], and ControlNet [60]. Finally, we ablate each component of our method and show its effectiveness.
Rows 1-3 of Fig. 5 present examples of images applied to three classic photographic effects: HDR, photo-tile-shift, and long-exposure. Rows 4-9 of Fig. 5 show the results of transfer a single image to various art styles. Our method successfully preserves the structure and natural appearance of the image, while enabling effective style transfer.
@@
Figure 6: Comparison of style transfer results using different methods. From left to right: the source content image and text prompt, our results, FreeDiff, P2P, PnP, and SDEdit with $0.6$ denoising strengths. Zoom in for the best view.
Figure 7: Comparison of visual enhancement results using different methods. From left to right: the source content image, visual enhancement prompts, our results, FreeDiff, P2P, PnP, and SDEdit with $0.6$ denoising strengths. Zoom in for the best view.
Figure 8: Qualitative comparisons with various style transfer methods. Left: Anime style transfer, where AdaIN, AdaAttn, and Z* are style-image-driven style transfer methods. right: Background blur for visual enhancement,which image-driven methods fail to achieve due to the lack of a reference image.
V-A Implementation Details
We conduct comparison experiments by loading the open-source Realistic-Vision-V4 weights [39] into the Stable Diffusion V1 model, which generates images closely resembling those in the real world. We use DDIM deterministic sampling with 50 steps in all of our experiments. In the case of real guidance images, we conduct deterministic DDIM inversion with $1,000$ forward steps, followed by deterministic DDIM sampling with 1000 backward steps to achieve a relative precise inversion. Our stylized results are performed with 50 sampling timesteps, thus we save the corresponding latents every 25 timesteps for use in each step of the 50 timesteps style transfer process. In addition, we set the classifier-free guidance scale for real and generated guidance images to 7.5. The ASI module only replaces the cross-attention module of the diffusion model in the conditional branch. For Stable Diffusion V1 model, we set the default thresholds for the spatial mask to $\alpha=0.7$, and set the attention head mask to $n=6$ and $n=7$ for visual enhancement and style transfer tasks, respectively. For Stable Diffusion XL model, we set the $\alpha=0.6$, $n=16$ and $n=18$ for visual enhancement and style transfer tasks, respectively. In addition, following the PNP [50], during the editing process, we replace the self-attention scores from layers 4 to 11 of the UNet with the corresponding self-attention scores from the reference diffusion process for the first 25 timesteps. Similarly, we also replace the convolution features from layer 4 of the UNet during the first 40 timesteps.
V-B Qualitative Comparison
V-B1 Style Transfer
To verify the effectiveness and generalization of ASI, we show various style transfer results applied by our approach for real-world images and generated images. For each result, we show images before and after stylized. As shown in Fig. 6, we compare our method with SOTA text-guided style transfer methods based on diffusion models, including FreeDiff [55], P2P [20], PNP [50], and SDEdit [33]. The first three rows and the last three rows respectively show the results of style transfer based on real and generated images. Since our method aims to incorporate style information at the fine-grained feature level, compared to prompt-level style injection, it allows us to stylize in the smaller local region of images, thereby reducing the distortion of image structure. Although SDEdit is simple but effective, it requires manual adjustment of the trade-off between the two. FreeDiff can preserve object consistency in stylized images, but it struggles to accurately maintain the image’s structural consistency (Row. 1,2,4-6) and effective stylization (Row. 3). PNP and P2P either have excessive stylization(e.g., Row. 1,2,4, and 5) leading to structural and instance deviation, or they lack effective stylization(Row. 3 and 4). As shown in Col. 2 of Fig. 6, our method effectively stylizes images while maintaining structural consistency with the source content image, achieving a good balance between stylization and structural preservation compared to other SOTA methods.
V-B2 Visual Enhancement